OpenStudy (anonymous):

to find general solution of (x^2-1)y"-2xy'+2y=(x^2-1)^2

OpenStudy (anonymous):

we have to solve it by method of variation of parameters

OpenStudy (anonymous):

Here is what I got from Mathematica \[ y(x)=\frac{c_1 \sqrt{x^2-1} (1-x)^{3/2}}{\sqrt{x+1}}-\\ \frac{c_2 x \sqrt{x^2-1} \sqrt{1-x}}{(x-1) \sqrt{x+1}}+\\ \frac{3 \sqrt{1-x} x^5-3 \sqrt{1-x} x^4-3 \sqrt{1-x} x^3+3 \sqrt{1-x} x^2-6 \sqrt{x+1} \sqrt{1-x^2} x^2+2 \sqrt{x+1} \sqrt{1-x^2} x-2 \sqrt{x+1} \sqrt{1-x^2} x^4+6 \sqrt{x+1} \sqrt{1-x^2} x^3}{6 \sqrt{x+1} \sqrt{1-x^2}} \]

OpenStudy (anonymous):

Here is another form \[ y(x)=\frac{6 c_1 \sqrt{-\left(x^2-1\right)^2} x^2}{6-6 x^2}+ \frac{6 c_2 \sqrt{-\left(x^2-1\right)^2} x}{6-6 x^2}-\\ \frac{12 c_1 \sqrt{-\left(x^2-1\right)^2} x}{6-6 x^2}+ \frac{6 c_1 \sqrt{-\left(x^2-1\right)^2}}{6-6 x^2}- \frac{3 x^2}{6-6 x^2}+\frac{2 x}{6-6 x^2}-\\ \frac{x^6}{6-6 x^2}+\frac{4 x^4}{6-6 x^2}-\frac{2 x^3}{6-6 x^2} \]

OpenStudy (anonymous):

This should be done y a machine and not by hand.